if sinx+sin2x=2


then find,cos2x+cos4x

Grade:10

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
sin^{2}x + sinx = 2
(sinx+2)(sinx-1) = 0
sinx = 1
is the only solution.
x = n\pi \pm \frac{\pi }{2}
cos2x + cos4x = cos2(n\pi \pm \frac{\pi }{2}) + cos4(n\pi \pm \frac{\pi }{2})
cos2x + cos4x = cos(2n\pi \pm \pi ) + cos(4n\pi \pm 2\pi ) = -1 + 1 = 0
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty